OpenStudy (anonymous):

A car dealer offers you two deals on a car that costs $16,000. Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $3,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 1% simple annual interest and no down payment. (Hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time (in years). Add the amount of interest to the price of the car.)

OpenStudy (anonymous):

16000-3000= 13000/60=216.67 month for 60 months

OpenStudy (anonymous):

opt 2 16000*.01*72=960+16000=16960/72=235.56 a month for 72 months

OpenStudy (anonymous):

16000-1%*6 (years not 72 months)

OpenStudy (anonymous):

16000*1% - sorry cant type tonight